### v1

# Quasi-two-body decays $B\to Pf_{0}(500)\to P\pi^{+}\pi^{-}$ in the perturbative QCD approach

Abstract
[…]
pacs:
13.20.He, 13.25.Hw, 13.30.Eg
I INTRODUCTION
Three-body $B$-meson decays are considerably more challenging than that of two-body-decays, mainly due to the entangled resonant and nonresonant contributions, the complex interplay between the weak and strong dynamics Charles et al. (2017), and other possible final-state-interactions (FSI) Bediaga and Magalhães (2015); Bediaga et al. (2018); Yu et al. (2024) in the three-body $B$ meson decays. Traditional approaches for the two-body-decays are no longer satisfactory in the three-body decay processes Alvarenga Nogueira et al. (2016). Practically, the hadronic three-body $B$ meson decay processes, in most cases, are considered to be dominated by the low-energy $S$-, $P$- and $D$-wave resonant states, which could be treated in the quasi-two-body framework. By neglecting the FSI between the meson pair originated from the resonant states and the bachelor particle, the factorization procedure can be applied Alvarenga Nogueira et al. (2016); Boito et al. (2017). Substantial theoretical efforts for different quasi-two-body $B$ meson decays has been made within different theoretical approaches, c.f. Refs. Gronau and Rosner (2003); Engelhard et al. (2005); Gronau and Rosner (2005); Imbeault and London (2011); Gronau (2013); Bhattacharya et al. (2013, 2014); Xu et al. (2014a, b); He et al. (2015); Furman et al. (2005); El-Bennich et al. (2006, 2009); Leitner et al. (2010); Dedonder et al. (2011); Cheng et al. (2005, 2007); Cheng and Chua (2013, 2014); Li (2014); Cheng et al. (2016); Kränkl et al. (2015); Klein et al. (2017). As well, the contributions from various intermediate resonant states for the three-body $B$-meson decays in the context of perturbative QCD (PQCD) approach Keum et al. (2001a, b); Lü et al. (2001) have been investigated in Ref. Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022).
Compared with vector and tensor mesons, the identification of the scalar mesons is long-standing puzzle Workman et al. (2022). Scalar resonances are hard to be resolved, since some of them have large decay widths which make us difficult to distinguish between resonance and background. In the theoretical point of view, their masses do not fit the expectation in the naive quark model. For the lightest scalar meson $f_{0}(500)$ (also refereed as $\sigma$) meson, which is, in the present statu, not a ordinary meson in the sense that it cannot be interpreted as predominantly made of quark and antiquark Pelaez (2016). After more than 60 years of study of $f_{0}(500)$, various of interpretations have been proposed. More explicitly, a light scalar-isoscalar field was first postulated Johnson and Teller (1955) for explanation of the inter-nucleon attraction. Then, Linear Sigma Model was proposed Gell-Mann and Levy (1960) to describe the chiral symmetry in pion-pion interaction, this explains why $f_{0}(500)$ is usually called as the $\sigma$ meson. The linear sigma model plays a relevant role, in history, for the understanding of spontaneous chiral symmetry breaking, where all fields become Goldstone bosons, i.e. pions, except $\sigma$. When the $f_{0}(500)$ and $f_{0}(980)$ are considered as $q\bar{q}$ state, the mixing of light and strange quark may appear Stone and Zhang (2013), it can be characterized by a $2\times 2$ rotation matrix with a single parameter, i.e. the mixing angle $\phi$. If the $q\bar{q}$ state does exist, the mixing angle can be constrained by the scalar decay channel which is exactly what we study. Comparison between our theoretical prediction and the experimental data may allow us to probe the inner structure of light scalar mesons. Tetraquarks is also a popular interpretation Jaffe (1977) which is bounded to form a color neutral resonance by two valence quark and two antiquark Pelaez (2016). There are also some different explanation of quark level dynamics to describe the formation of scalar mesons, such as ‘bag model’ with additional one gluon exchange Jaffe (1977), the diquark-antidiquark configurations Maiani et al. (2004), the large-$N$ quantum chromodynamics Weinberg (2013), moreover, including instanton effects ’t Hooft et al. (2008). In this theories, tetraquark states are unmixed, as assumed in Ref. Stone and Zhang (2013), with a constrain of mixing angle less than 5 degree Fleischer et al. (2011). Consequently, how to distinguish $q\bar{q}$ and $q\bar{q}q\bar{q}$ states becomes a significant issue. In Ref. Wang and Lu (2010), authors proposed a method to distinguish two kinds of scalar mesons based on sum rule technique, it may becomes a important criterion of phenomenological study and may provide more detail of inner structure of scalar mesons.
Figure 1: Typical diagrams for the quasi-two-body decays $B_{(s)}\to P\sigma\to P\pi\pi$. The diagram (a) for the $B\to\sigma$ transition, and diagram (c) for the $B\to P$ transition, as well as the diagrams (b) and (d) for for annihilation contributions. The symbol $\otimes$ stands for the weak vertex and $\times$ denotes possible attachments of hard gluons.
Within the frame work of PQCD approach, quasi-two body decays involving resonant state $f_{0}(500)$ have been studied Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), we would like to extend to our previous studies to the quasi-two-body $B$ meson decays $B\to Pf_{0}(500)\to P\pi\pi$, where the bachelor particle $P$ denotes the light pseudoscalar $\pi$, $K$, $\eta$, or $\eta^{\prime}$. Typical diagrams for the $B\to Pf_{0}(500)\to P\pi\pi$ decays’ processes are shown in Fig. 1. Inspired by the generalized parton distribution (GPD) in hard exclusive two pion production Diehl et al. (1998); Müller et al. (1994); Polyakov (1999); Hägler et al. (2002), the two-meson distribution amplitude was introduced in three-body hadronic $B$ decays in the frame work of PQCD approach Chen and Li (2003, 2004) as the universal nonperturbative input. The decay amplitude for the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi\pi$ can be expressed as the convolution of the nonperturbative wave function and hard kernel Chen and Li (2003, 2004); Wang and Li (2016)
$\mathcal{A}=\phi_{B}\otimes H\otimes\phi_{P}\otimes\phi_{\pi\pi}^{S\text{-wave}},$ (1)
where hard kernel $H$ contains one hard gluon, and the distribution amplitudes $\phi_{B},\phi_{P}$ and $\phi_{\pi\pi}^{S\text{-wave}}$ absorb the nonperturbative dynamics in the decay processes.
The rest of this paper are organised as follows. We give a brief introduction of the theoretical framework in Sec. II. Numerical results and some discussions are shown in Sec. III, and a brief conclusion will be summarized in Sec. IV. The relevant factorization formulas for the decay amplitudes are collected in the Appendix.
II FRAMEWORK
In the light-cone coordinate, the $B$ meson momentum $p_{B}$, $\sigma$ meson momentum $p$, bachelor particle momentum $p_{3}$, and the corresponding quark momenta $k_{B}$, $k$, $k_{3}$ in the rest frame of $B$ meson are defined as
$p_{B}=\frac{m_{B}}{\sqrt{2}}(1,1,\textbf{0}_{\rm T}),$ $k_{B}=\left(0,\frac{m_{B}}{\sqrt{2}}x_{B},\textbf{k}_{B{\rm T}}\right),$
$p=\frac{m_{B}}{\sqrt{2}}(1,\eta,\textbf{0}_{\rm T}),$ $k=\left(\frac{m_{B}}{\sqrt{2}}z,0,\textbf{k}_{\rm T}\right),$
$p_{3}=\frac{m_{B}}{\sqrt{2}}(0,1-\eta,\textbf{0}_{\rm T}),$ $k_{3}=\left(0,\frac{m_{B}}{\sqrt{2}}(1-\eta)x_{3},\textbf{k}_{3{\rm T}}\right),$ (2)
where the variable $\eta$ is defined as $\eta=\omega^{2}/m_{B}^{2}$ with $\omega=\sqrt{p^{2}}$ stand for invariant mass of dipion. Terms $x_{B}$, $z$, and $x_{3}$ are the momentum fraction of $k_{B}$, $k$, and $k_{3}$ respectively.
The $B$ meson can be treated as heavy-light system, whose wave function can be written as Li (2003)
$\Phi_{B}(x,b)=\frac{1}{\sqrt{2N_{c}}}\left({p_{B}}/\ +m_{B}\right)\gamma_{5}\phi_{B}(x,b),$ (3)
where $b$ is the conjugate space coordinate of the transverse momentum of the valence quark of meson, $N_{c}$ is the color factor, the distribution amplitude $\phi_{B}(x,b)$ are chosen as
$\phi_{B}(x,b)=N_{B}x^{2}(1-x^{2})\exp\left[-\frac{1}{2}\left(\frac{xm_{B}}{\omega_{B}}\right)^{2}-\frac{\omega_{B}^{2}b^{2}}{2}\right],$ (4)
where the $\omega_{B}$ is the shape parameter and $N_{B}$ being the normalization factor. The shape parameter $\omega_{B}$ is mainly fixed from the fit to the $B\to\pi$ form factors derived from lattice QCD Bowler et al. (2000) and LCSR Ball (1998). The normalization constant $N_{B}$ is related to the decay constant $f_{B}$ through the relation
[…]
For the final state of light pseudoscalar meson, whose wave function being the
$\Phi_{P}(p,z)=\frac{i}{\sqrt{2N_{c}}}\gamma_{5}\left[p/\phi^{A}(z)+m_{0}\phi^{P}(z)+m_{0}(v/n/-1)\phi^{T}(z)\right],$ (6)
where $m_{0}$ is chiral mass of corresponding pseudoscalar meson. The $p$ and $z$ are associated momentum and momentum fraction respectively. The explicit expression of relevant distribution amplitude in light-cone sum rule up to twist-3 are shown Ball (1998, 1999); Ball and Zwicky (2005); Ball et al. (2006)
$\phi^{A}(x)$ $=$ $\frac{f_{P}}{2\sqrt{2N_{c}}}6x(1-x)\left[1+a_{1}^{P}C_{1}^{3/2}(2x-1)+a_{2}^{P}C_{2}^{3/2}(2x-1)+a_{4}^{P}C_{4}^{3/2}(2x-1)\right],$
$\phi^{P}(x)$ $=$ $\frac{f_{P}}{2\sqrt{2N_{c}}}\bigg{[}1+\left(30\eta_{3}-\frac{5}{2}\rho^{2}_{P}\right)C_{2}^{1/2}(2x-1)-3\left[\eta_{3}\omega_{3}+\frac{9}{20}\rho^{2}_{P}(1+6a_{2}^{P})\right]C_{4}^{1/2}(2x-1)\bigg{]},$
$\phi^{T}(x)$ $=$ $\frac{f_{P}}{2\sqrt{2N_{c}}}(1-2x)\bigg{[}1+6\left(5\eta_{3}-\frac{1}{2}\eta_{3}\omega_{3}-\frac{7}{20}\rho^{2}_{P}-\frac{3}{5}\rho^{2}_{P}a_{2}^{P}\right)(1-10x+10x^{2})\bigg{]},$ (7)
with the Gegenbauer moments are $a_{1}^{\pi,\eta_{q,s}}=0,\ a_{1}^{K}=0.06,\ a_{2}^{\pi,K}=0.25,\ a_{2}^{\eta_{q,s}}=0.115,\ a_{4}^{\pi,\eta_{q,s}}=-0.015$. Meanwhile, the parameters are $\rho_{\pi}=m_{\pi}/m_{0}^{\pi},\ \rho_{K}=m_{K}/m_{0}^{K},\ \rho_{\eta_{q}}=2m_{q}/m_{0}^{q},\ \rho_{\eta_{s}}=2m_{s}/m_{0}^{s}$, $\eta_{3}=0.015,\ \omega_{3}=-3$ with $m_{0}^{\pi}=(1.4\pm 0.1)$ GeV, $m_{0}^{K}=(1.6\pm 0.1)$ GeV, $m_{0}^{\eta_{q}}=1.07$ GeV, $m_{0}^{\eta_{s}}=1.92$ GeV. The definition of Gegenbauer polynomials can be found in Ref. Ball and Zwicky (2005); Ball et al. (2006).
For the $\eta$ and $\eta^{\prime}$ mesons, its components has been studied extensively, two major mixing mechanism were adopt in most study, the quark flavor basis and singlet-octet basis. Different process under different assumption are studied to determine the mixing angle, we prefer to choose the so-called Feldmann-Kroll-Stech (FKS) formalism Feldmann et al. (1998, 1999) in which it was considered as mixing of $\eta_{q}$ and $\eta_{s}$, which made of $n\bar{n}=(u\bar{u}+d\bar{d})/\sqrt{2}$ and $s\bar{s}$ respectively, the physical state $\eta$ and $\eta^{\prime}$ related to flavor state $\eta_{q}$ and $\eta_{s}$ through a two by two rotation matrix with a single parameter, the mixing angle $\phi$
$\left(\begin{array}[]{c}|\eta\rangle\\ |\eta^{\prime}\rangle\end{array}\right)=\left(\begin{array}[]{cc}\text{cos}\phi&-\text{sin}\phi\\ \text{sin}\phi&\text{cos}\phi\end{array}\right)\left(\begin{array}[]{c}|\eta_{q}\rangle\\ |\eta_{s}\rangle\end{array}\right)$ (14)
with $\phi=39.3^{\circ}\pm 1.0^{\circ}$. For the possible glueball effect are considered to be small Charng et al. (2006), we will neglect this contribution. The distribution amplitude of $n\bar{n}$ and $s\bar{s}$ contain same Lorentz structure with pion except for the difference decay constant and chiral parameters, we collect the relation Charng et al. (2006)
$f_{n}$ $=$ $(1.07\pm 0.02)f_{\pi}=139.1\pm 2.6~{}\text{MeV},$
$f_{s}$ $=$ $(1.34\pm 0.06)f_{\pi}=174.2\pm 7.8~{}\text{MeV}.$ (15)
The $S$-wave $\pi$-$\pi$ distribution amplitude takes the following form of Diehl et al. (1998); Müller et al. (1994); Polyakov (1999); Hägler et al. (2002); Meißner and Wang (2014)
$\Phi_{\pi\pi}^{S\text{-wave}}=\frac{1}{\sqrt{2N_{c}}}\bigg{[}p/\phi^{I=0}_{v\nu=-}(z,\zeta,\omega^{2})+\omega\phi^{I=0}_{s}(z,\zeta,\omega^{2})+\omega(n/v/-1)\phi^{I=0}_{t\nu=+}(z,\zeta,\omega^{2})\bigg{]},$ (16)
where
$\phi^{I=0}_{v\nu=-}(z,\zeta,\omega^{2})$ $=$ $\phi^{0}=\frac{9F_{s}(\omega^{2})}{\sqrt{2N_{c}}}a_{2}^{I=0}z(1-z)(1-2z),$
$\phi^{I=0}_{s}(z,\zeta,\omega^{2})$ $=$ $\phi^{s}=\frac{F_{s}(\omega^{2})}{2\sqrt{2N_{c}}},$
$\phi^{I=0}_{t\nu=+}(z,\zeta,\omega^{2})$ $=$ $\phi^{t}=\frac{F_{s}(\omega^{2})}{2\sqrt{2N_{c}}}(1-2z),$ (17)
where the expression of scalar form factor and associated auxiliary functions can be found in Wang et al. (2015); Wang (2019).
According to the typical Feynman diagrams as shown in Fig. 1 and the quark currents for each decays, the decay amplitudes for considered quasi-two-body decays $B\to Pf_{0}(500)\to P\pi\pi$ are given as
${\mathcal{A}}\big{(}B^{+}\to\pi^{+}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}V^{*}_{ub}V_{ud}\bigg{[}\left(\frac{C_{1}}{3}+C_{2}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}+F^{LL}_{AP}\right)+C_{1}\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}+M^{LL}_{AP}\right)$ (18)
$+$ $C_{2}M^{LL}_{TP}\bigg{]}-V^{*}_{tb}V_{td}\bigg{[}\left(\frac{C_{3}}{3}+C_{4}+\frac{C_{9}}{3}+C_{10}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}+F^{LL}_{AP}\right)$
$+$ $\left(\frac{C_{5}}{3}+C_{6}+\frac{C_{7}}{3}+C_{8}\right)\left(F^{SP}_{T\sigma}+F^{SP}_{A\sigma}+F^{SP}_{AP}\right)$
$+$ $\left(C_{3}+C_{9}\right)\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}+M^{LL}_{AP}\right)+\left(C_{5}+C_{7}\right)\left(M^{LR}_{T\sigma}+M^{LR}_{A\sigma}+M^{LR}_{AP}\right)$
$+$ $\left(C_{3}+2C_{4}-\frac{C_{9}}{2}+\frac{C_{10}}{2}\right)M^{LL}_{TP}\left(C_{5}-\frac{C_{7}}{2}\right)M^{LR}_{TP}$
$+$ $\left(2C_{6}+\frac{C_{8}}{2}\right)M^{SP}_{TP}+\left(\frac{C_{5}}{3}+C_{6}-\frac{C_{7}}{6}-\frac{C_{8}}{2}\right)F_{TP}^{SP}\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{+}\to K^{+}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}V^{*}_{ub}V_{us}\bigg{[}\left(\frac{C_{1}}{3}+C_{2}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}\right)+C_{1}\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}\right)+C_{2}M^{LL}_{TP}\bigg{]}$ (19)
-- $V^{*}_{tb}V_{ts}\bigg{[}\left(\frac{C_{3}}{3}+C_{4}+\frac{C_{9}}{3}+C_{10}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}\right)$
$+$ $\left(\frac{C_{5}}{3}+C_{6}+\frac{C_{7}}{3}+C_{8}\right)\left(F^{SP}_{T\sigma}+F^{SP}_{A\sigma}\right)+\left(C_{3}+C_{9}\right)\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}\right)$
$+$ $\left(C_{5}+C_{7}\right)\left(M^{LR}_{T\sigma}+M^{LR}_{A\sigma}\right)+\left(2C_{4}+\frac{C_{10}}{2}\right)M^{LL}_{TP}+\left(2C_{6}+\frac{C_{8}}{2}\right)M^{SP}_{TP}\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{0}\to\pi^{0}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2\sqrt{2}}\bigg{\{}V^{*}_{ub}V_{ud}\bigg{[}\left(C_{1}+\frac{C_{2}}{3}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}+F^{LL}_{AP}\right)$ (20)
$+$ $C_{2}\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}+M^{LL}_{TP}+M^{LL}_{AP}\right)\bigg{]}$
-- $V^{*}_{tb}V_{td}\bigg{[}\left(-\frac{C_{3}}{3}-C_{4}-\frac{3C_{7}}{2}-\frac{C_{8}}{2}+\frac{5C_{9}}{3}+C_{10}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}+F^{LL}_{AP}\right)$
$+$ (−$\left(-\frac{C_{5}}{3}-C_{6}+\frac{C_{7}}{6}+\frac{C_{8}}{2}\right)\left(F^{SP}_{T\sigma}+F^{SP}_{A\sigma}+F^{SP}_{TP}+F^{SP}_{AP}\right)$
$+$ (−$\left(-C_{3}+\frac{C_{9}}{2}+\frac{3C_{10}}{2}\right)\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}+M^{LL}_{AP}\right)$
$+$ (−$\left(-C_{3}-2C_{4}+\frac{C_{9}}{2}-\frac{C_{10}}{2}\right)M_{TP}^{LL}+\frac{3C_{8}}{2}\left(M^{SP}_{T\sigma}+M^{SP}_{A\sigma}+M^{SP}_{AP}\right)$
$+$ (−$\left(-2C_{6}-\frac{C_{8}}{2}\right)M^{SP}_{TP}+\left(-C_{5}+\frac{C_{7}}{2}\right)\left(M^{LR}_{T\sigma}+M^{LR}_{A\sigma}+M^{LR}_{TP}+M^{LR}_{AP}\right)\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{0}\to K^{0}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}V^{*}_{ub}V_{us}\bigg{[}C_{2}M^{LL}_{TP}\bigg{]}-V^{*}_{tb}V_{ts}\bigg{[}\left(\frac{C_{3}}{3}+C_{4}-\frac{C_{9}}{6}-\frac{C_{10}}{2}\right)\left(F^{LL}_{T\sigma}+F^{LL}_{A\sigma}\right)$ (21)
$+$ $\left(\frac{C_{5}}{3}+C_{6}-\frac{C_{7}}{6}-\frac{C_{8}}{2}\right)\left(F^{SP}_{T\sigma}+F^{SP}_{A\sigma}\right)+\left(C_{3}-\frac{C_{9}}{2}\right)\left(M^{LL}_{T\sigma}+M^{LL}_{A\sigma}\right)$
$+$ $\left(C_{5}-\frac{C_{7}}{2}\right)\left(M^{LR}_{T\sigma}+M^{LR}_{A\sigma}\right)+\left(2C_{4}+\frac{C_{10}}{2}\right)M^{LL}_{TP}+\left(2C_{6}+\frac{C_{8}}{2}\right)M^{SP}_{TP}\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B_{s}^{0}\to\bar{K}^{0}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}V^{*}_{ub}V_{ud}\bigg{[}C_{2}M_{TP}^{LL}\bigg{]}-V^{*}_{tb}V_{td}\bigg{[}\left(\frac{C_{5}}{3}+C_{6}-\frac{C_{7}}{6}-\frac{C_{8}}{2}\right)\left(F^{SP}_{TP}+F^{SP}_{AP}\right)$ (22)
$+$ $\left(\frac{C_{3}}{3}+C_{4}-\frac{C_{9}}{6}-\frac{C_{10}}{2}\right)F_{AP}^{LL}+\left(C_{3}+2C_{4}-\frac{C_{9}}{2}+\frac{C_{10}}{2}\right)M_{TP}^{LL}$
$+$ $\left(C_{5}-\frac{C_{7}}{2}\right)M_{TP}^{LR}+\left(2C_{6}-\frac{C_{8}}{2}\right)M_{TP}^{SP}+\left(C_{3}-\frac{C_{9}}{2}\right)M_{AP}^{LL}$
$+$ $\left(C_{5}-\frac{C_{7}}{2}\right)M_{AP}^{LR}\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2\sqrt{2}}\bigg{\{}V^{*}_{ub}V_{ud}\bigg{[}\left(C_{1}+\frac{C_{2}}{3}\right)\left(F_{T\sigma}^{LL}+F_{A\sigma}^{LL}+F_{AP}^{LL}\right)$ (23)
$+$ $C_{2}\left(M_{T\sigma}^{LL}+M_{A\sigma}^{LL}+M_{TP}^{LL}+M_{AP}^{LL}\right)\bigg{]}$
-- $V^{*}_{tb}V_{td}\bigg{[}\left(C_{5}+\frac{C_{6}}{3}-\frac{C_{7}}{6}-\frac{C_{8}}{2}\right)\left(F^{SP}_{T\sigma}+F^{SP}_{TP}\right)+\left(C_{6}-\frac{C_{8}}{2}\right)\left(M_{T\sigma}^{SP}+M_{TP}^{SP}\right)$
$+$ $\left(\frac{7C_{3}}{3}+\frac{7C_{4}}{3}-2C_{5}-\frac{2C_{6}}{3}-\frac{C_{7}}{2}-\frac{C_{8}}{6}+\frac{C_{9}}{3}-\frac{C_{10}}{3}\right)\left(F_{T\sigma}^{LL}+F_{A\sigma}^{LL}+F_{AP}^{LL}\right)$
$+$ $\left(C_{3}+2C_{4}-\frac{C_{9}}{2}+\frac{C_{10}}{2}\right)\left(M_{T\sigma}^{LL}+M_{TP}^{LL}\right)$
$+$ $\left(C_{5}+C_{6}-\frac{C_{7}}{2}+C_{8}\right)\left(M_{T\sigma}^{LR}+M_{TP}^{LR}\right)$
$+$ $\left(\frac{C_{5}}{3}+C_{6}-\frac{C_{7}}{6}-\frac{C_{8}}{2}\right)\left(F_{A\sigma}^{SP}+M_{AP}^{SP}\right)+\left(C_{4}+C_{10}\right)\left(M_{A\sigma}^{LL}+M_{AP}^{LL}\right)$
$+$ $\left(C_{6}+C_{8}\right)\left(M_{A\sigma}^{SP}+M_{AP}^{SP}\right)\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}-V^{*}_{tb}V_{td}\bigg{[}\left(C_{4}-\frac{C_{10}}{2}\right)M_{T\sigma}^{LL}+\left(C_{6}-\frac{C_{8}}{2}\right)M_{T\sigma}^{SP}$ (24)
$+$ $\left(C_{3}+\frac{C_{4}}{3}-C_{5}-\frac{C_{6}}{3}+\frac{C_{7}}{2}+\frac{C_{8}}{6}-\frac{C_{9}}{2}-\frac{C_{10}}{6}\right)F_{T\sigma}^{LL}\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B^{0}\to\eta[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ ${\mathcal{A}}\big{(}B^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}\cos\phi-{\mathcal{A}}\big{(}B^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}\sin\phi\;,$ (25)
${\mathcal{A}}\big{(}B^{0}\to\eta^{\prime}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ ${\mathcal{A}}\big{(}B^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}\sin\phi+{\mathcal{A}}\big{(}B^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}\cos\phi\;,$ (26)
${\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2\sqrt{2}}\bigg{\{}V^{*}_{ub}V_{us}\bigg{[}\left(C_{1}+\frac{C_{2}}{3}\right)\left(F_{A\sigma}^{LL}+F_{AP}^{LL}\right)+C_{2}\left(M_{A\sigma}^{LL}+M_{AP}^{LL}\right)\bigg{]}$ (27)
-- $V^{*}_{tb}V_{ts}\bigg{[}\left(2C_{3}+\frac{2C_{4}}{3}-2C_{5}-\frac{2C_{6}}{3}-\frac{C_{7}}{2}-\frac{C_{8}}{6}+\frac{C_{9}}{2}+\frac{C_{10}}{6}\right)\left(F_{A\sigma}^{LL}+F_{AP}^{LL}\right)$
$+$ $\left(2C_{4}+\frac{C_{10}}{2}\right)\left(M_{A\sigma}^{LL}+M_{AP}^{LL}\right)+\left(2C_{6}+\frac{C_{8}}{2}\right)\left(M_{A\sigma}^{SP}+M_{AP}^{SP}\right)\bigg{]}\bigg{\}}\;,$
${\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ $\frac{G_{F}}{2}\bigg{\{}V^{*}_{ub}V_{us}\bigg{[}C_{2}M_{TP}^{LL}\bigg{]}-V^{*}_{tb}V_{ts}\bigg{[}\left(C_{4}+C_{10}\right)M_{TP}^{LL}+\left(C_{6}+C_{8}\right)M_{TP}^{SP}\bigg{]}\bigg{\}}\;,$ (28)
${\mathcal{A}}\big{(}B_{s}^{0}\to\eta[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ ${\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}\cos\phi-{\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}\sin\phi\;,$ (29)
${\mathcal{A}}\big{(}B_{s}^{0}\to\eta^{\prime}[\sigma\to]\pi^{+}\pi^{-}\big{)}$ $=$ ${\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{q}[\sigma\to]\pi^{+}\pi^{-}\big{)}\sin\phi+{\mathcal{A}}\big{(}B_{s}^{0}\to\eta_{s}[\sigma\to]\pi^{+}\pi^{-}\big{)}\cos\phi\;.$ (30)
where $G_{F}$ is the Fermi constant, $V_{ij}$ is the CKM matrix element, and the combinations of the Wilson coefficients $a_{1,2}$ are defined as $a_{1}=C_{1}/3+C_{2}$ and $a_{2}=C_{2}/{3}+C_{1}$. The expressions of individual amplitudes $F^{LL}_{T\sigma}$, $F^{SP}_{T\sigma}$, $F^{LL}_{A\sigma}$, $F^{SP}_{A\sigma}$, $M^{LL}_{T\sigma}$, $M^{LR}_{T\sigma}$, $M^{SP}_{T\sigma}$, $M^{LL}_{A\sigma}$, $M^{LR}_{A\sigma}$, $M^{SP}_{A\sigma}$, $F^{SP}_{TP}$, $M^{LL}_{TP}$, $M^{LR}_{TP}$, $M^{SP}_{TP}$, $F^{LL}_{AP}$, $F^{SP}_{AP}$, $M^{LL}_{AP}$, $M^{LR}_{AP}$ and $M^{SP}_{AP}$ from different subdiagrams in Fig. 1, which are collected in the Appendix A.
Differential branching ratio of $B\to Pf_{0}(500)\to P\pi\pi$ can be written as
[…]
where $\tau_{B}$ is the mean lifetime of $B$ meson, $|\vec{p}_{1}|$ and $|\vec{p}|$ are the three momenta of $f_{0}(500)$ resonance and $D$ meson respectively in the center-of-mass frame of $\pi$-$\pi$, and can be written as
$|\overrightarrow{p_{1}}|=\frac{\sqrt{\lambda(\omega^{2},m_{\pi}^{2},m^{2}_{\pi})}}{2\omega},\quad|\overrightarrow{p}|=\frac{\sqrt{\lambda(m_{B}^{2},m_{D}^{2},\omega^{2})}}{2\omega},$ (32)
with $m_{B}$, $m_{D}$ and $m_{\pi}$ are the masses of the $B$, $D$ and $\pi$ mesons respectively, and the Källén function $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$.
III RESULTS
Parameter Value Unit Reference Parameter Value Unit Reference
$m_{B^{\pm}}$ 5.279 GeV Workman et al. (2022) $\tau_{B^{\pm}}$ 1.638 ps Workman et al. (2022)
$m_{B^{0}}$ 5.280 GeV Workman et al. (2022) $\tau_{B^{0}}$ 1.519 ps Workman et al. (2022)
$m_{B_{s}}$ 5.367 GeV Workman et al. (2022) $\tau_{B_{s}}$ 1.527 ps Workman et al. (2022)
$f_{B}|_{N_{f}=2+1+1}$ 190.0 MeV Aoki et al. (2022) $f_{B_{s}}|_{N_{f}=2+1+1}$ 230.3 MeV Aoki et al. (2022)
$m_{\pi^{\pm}}$ 0.140 GeV Workman et al. (2022) $m_{\pi^{0}}$ 0.135 GeV Workman et al. (2022)
$m_{K^{\pm}}$ 0.494 GeV Workman et al. (2022) $m_{K^{0}}$ 0.498 GeV Workman et al. (2022)
$m_{\eta}$ 0.548 GeV Workman et al. (2022) $m_{\eta^{\prime}}$ 0.958 GeV Workman et al. (2022)
$f_{\pi}$ 0.130 GeV Workman et al. (2022) $f_{K}$ 0.156 GeV Workman et al. (2022)
$\lambda$ 0.2250 Workman et al. (2022) $\bar{\rho}$ 0.159 Workman et al. (2022)
$A$ 0.826 Workman et al. (2022) $\bar{\eta}$ 0.348 Workman et al. (2022)
Table 1: Numerical values of the theory input parameters employed in the PQCD predictions of the quasi-two-body decays →$\to Pf(500)\to P\pi\pi$ as well as the subsequent phenomenological analysis for the quasi-two-body decay observables.
For the numerical calculation, we adopt QCD scale at $\mu=0.25$ GeV in modified minimal subtraction scheme. The decay constant of $B^{0,\pm}$, $B_{s}$ and light pseudoscalar mesons come from FLAG working group’s result Aoki et al. (2022). The masses and mean life times of $B^{0,\pm}$ and $B_{s}$ mesons, the masses of light pseudoscalar mesons and Wolfenstein parameters are all come from recent updated Review of Particle Physics Workman et al. (2022). We summarize explicitly the numerical values of the necessary standard model inputs and the hadronic parameters in Table 1.
By using the formula of differential branching fraction in Eq. (31) and explicit decay amplitude in Appendix A, we obtain the PQCD predictions of CP averaged branching fractions and direct CP violations in Table 2 for quasi-two-body $B\to Pf(500)\to P\pi\pi$ processes. The first uncertainty comes from shape parameter $\omega_{B}$ in the $B$ meson distribution amplitude, we vary it value at 10% magnitude, that is $\omega_{B^{0,\pm}}=0.40\pm 0.04$ GeV and $\omega_{B_{s}}=0.5\pm 0.05$ GeV for $B^{0,\pm}$ and $B_{s}$ mesons respectively. The second uncertainty comes from the Gegenbaur moment $a_{2}^{I=0}=0.20\pm 0.20$ in the $S$-wave $\pi$-$\pi$ distribution amplitude. We ignore the uncertainties of parameters in distribution amplitude of light pseudoscalar mesons and Wolfenstein parameters since this uncertainties are considered to be small. It is interesting to see that the $\omega_{B}$ produce the largest theoretical uncertainty in PQCD predictions of quasi-two-body $B$ meson decays Wang and Li (2016); Ma et al. (2017a, b); Rui and Wang (2018); Li et al. (2018); Wang and Chai (2019); Ma et al. (2019); Cui et al. (2019); Wang et al. (2020); Fan and Wang (2020); Chai et al. (2022), but in this study and previous quasi-two-body $B$ meson decays involving $S$-wave $\pi$-$\pi$ contributions Wang et al. (2015); Ma et al. (2017a); Liu et al. (2019); Xing and Xing (2019), the Gegenbaur moment $a_{2}^{I=0}$ plays more important role.
Decay modes    Quasi-two-body results
$B^{+}\to\pi^{+}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-6})$ $1.78\pm 0.41\;(\omega_{B})\pm 0.51\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$ $29.8\pm 11.1\;(\omega_{B})\pm 13.0\;(a_{2})$
$B^{+}\to K^{+}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-7})$ $8.14\pm 1.26\;(\omega_{B})\pm 1.10\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$$-63.5\pm 8.2\;(\omega_{B})\pm 11.3\;(a_{2})$
$B^{0}\to\pi^{0}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-7})$ $2.02\pm 0.55\;(\omega_{B})\pm 1.23\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$$-64.6\pm 21.8\;(\omega_{B})\pm 38.7\;(a_{2})$
$B^{0}\to K^{0}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-7})$ $3.90\pm 0.26\;(\omega_{B})\pm 2.68\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$ $1.31\pm 18.0\;(\omega_{B})\pm 10.6\;(a_{2})$
$B_{s}^{0}\to\bar{K}^{0}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-7})$ $1.18\pm 0.22\;(\omega_{B})\pm 0.72\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$ $2.26\pm 8.61\;(\omega_{B})\pm 8.13\;(a_{2})$
$B^{0}\to\eta[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-8})$ $5.11\pm 1.11\;(\omega_{B})\pm 4.60\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$$-88.9\pm 1.3\;(\omega_{B})\pm 47.24\;(a_{2})$
$B^{0}\to\eta^{\prime}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-8})$ $2.29\pm 0.61\;(\omega_{B})\pm 1.59\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$$-42.3\pm 13.4\;(\omega_{B})\pm 5.4\;(a_{2})$
$B_{s}^{0}\to\eta[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-9})$ $2.00\pm 1.04\;(\omega_{B})\pm 1.25\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$ $15.2\pm 35.1\;(\omega_{B})\pm 17.1\;(a_{2})$
$B_{s}^{0}\to\eta^{\prime}[\sigma\to]\pi^{+}\pi^{-}$ ℬ$\mathcal{B}(10^{-8})$ $2.33\pm 0.72\;(\omega_{B})\pm 1.05\;(a_{2})$
$\mathcal{A}_{\rm CP}(\%)$ $10.4\pm 3.9\;(\omega_{B})\pm 6.8\;(a_{2})$
Table 2: PQCD predictions of CP-averaged branching fraction and direct CP violation for the quasi-two-body decays $B\to Pf(500)\to P\pi\pi$.
From the numerical results as listed in Table 2, we have the following comments:
    In the $B\to PR\to P\pi\pi$ decays, we can extract the two-body branching fractions ℬ$\mathcal{B}(B\to PR)$ by using the relation under the quasi-two-body approximation
[…]
    Combined with results listed in Table 2, one can obtain the related two-body branching fractions for the two-body decays $B\to Pf_{0}(500)$. It is interesting to see that the finite width effect are prominent in $B\to Pf_{0}(500)$ decay Cheng et al. (2021), which means the extraction of two-body results by using quasi-two-body approximation will be greatly underestimated.
    The predicted CP-averaged branching fractions for the considered decay processes are in the range of $10^{-9}-10^{-6}$. For the $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ decay, our prediction $(1.78\pm 0.41\pm 0.51)\times 10^{-6}$ is agree with Ref. Cheng et al. (2021) from the QCDF prediction by considering the finite width effect, both of two predictions satisfy the upper limit $4.1\times 10^{-6}$ from BABAR measurement Aubert et al. (2005). However, most resent LHCb measurement Aaij et al. (2020a, b) for the $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ decay in the context of isobar model yields $(3.83\pm 0.84)\times 10^{-6}$, the PQCD predicted branching fraction is smaller than the LHCb measurement by a factor of about 2. To have a clear look at the above comparison, we present the predictions in Table 3. Furthermore, we show the curve of $\pi$-$\pi$ invariant mass dependent differential branching fraction for quasi-two-body decay $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ in Fig. 2. It is found that the main portion of branching fraction for $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ received from the region of $[2\,m_{\pi},1\,{\rm GeV}]$, the contributions from the $m_{\pi\pi}>2.0$GeV is evaluated about $2.4\%$ and can be neglected safely.
    References       ℬ$\mathcal{B}(10^{-6})$       $\mathcal{A}_{\rm CP}(\%)$
    This Work 1.78Âą0.41Âą0.51plus-or-minus1.780.410.511.78$\pm$ 0.41$\pm$ 0.511.78 Âą 0.41 Âą 0.51 29.8Âą11.1Âą13.0plus-or-minus29.811.113.029.8$\pm$ 11.1$\pm$ 13.029.8 Âą 11.1 Âą 13.0
    QCDF Cheng et al. (2021) $1.65^{+0.42}_{-0.37}$ $14.7\pm 0.1$
    BABAR Aubert et al. (2005) $<4.1$
    LHCb Aaij et al. (2020a, b) $3.83\pm 0.84$ $14.9^{+0.5}_{-0.6}$
    Table 3: Predictions about CP-averaged branching fraction and direct CP violation for the quasi-two-body decays $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ decay. Meanwhile, we also listed the QCDF, BABAR, LHCb results as a comparison.
    Figure 2: The $\pi\pi$ invariant mass-dependent differential branching fraction for quasi-two-body decay $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$, the uncertainties generated from $\omega_{B}$ and $a_{2}$ are shown by the shaded band.
    For the quasi-two-body decay $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$, our prediction of CP violation reads $(29.8\pm 11.1\pm 13.0)\%$, which is much greater than that $14.9^{+0.5}_{-0.6}\%$ from LHCb measurements Aaij et al. (2020a, b) in the context of isobar model for the $S$-wave $\pi$-$\pi$, fortunately, this result is still in the range of our prediction by considering the uncertainty from two nonperturbative parameters, which can also be seen in Table 3. We also mention that QCDF prediction Cheng et al. (2021) is in excellent agreement with LHCb result. Besides, LHCb collaboration found that the interference between the $S$- and $P$-waves can also generate CP violation effect, unfortunately, CP violation can only produced from interference of tree and penguin diagrams in the state of art of PQCD calculations, the $S$- and $P$-waves cannot generate extra strong phase, the predictions of CP violation from interference of $S$- and $P$-waves is still absent and very challenging in PQCD. For sake of illustrate $\pi$-$\pi$ invariant mass dependent CP asymmetry, we show the $\pi$-$\pi$ invariant mass-dependent CP asymmetry for quasi-two-body decay $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ as an example, it is interesting to see that the direct CP asymmetry is decrease as the $\pi$-$\pi$ invariant mass increases.
    Figure 3: The $\pi\pi$ invariant mass-dependent CP asymmetry for quasi-two-body decay $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$, the uncertainties generated from $\omega_{B}$ and $a_{2}$ are shown by the shaded band.
    $f_{0}(500)$ was often parameterized by Bugg model Bugg (2007) in partial wave analysis and it was applied in LHCb measurements Aaij et al. (2015a, b). However, $S$-wave $\pi$-$\pi$ are considered to be broad, overlapping resonances, $K$-matrix model was applied for parameterization of $S$-wave components as a alternative scheme Aaij et al. (2020a, b). The rigorous theoretical calculation for nonresonance contribution in the context of PQCD framework is still absent Wang et al. (2020), comparison between experiment measurements and theoretical predictions is still challenging. More attempts can be make in future study to parameterize the nonresonance contribution for sake of giving a more reliable result.
IV CONCLUSION
We have studied the quasi-two-body decays $B\to Pf_{0}(500)\to P\pi^{+}\pi^{-}$ in the framework of PQCD factorization approach by using the scalar form factor $F_{\pi\pi}(\omega^{2})$ as the nonperturbative input. Here the bachelor particle $P$ denotes $\pi$, $K$, $\eta$ and $\eta^{\prime}$, respectively. The CP averaged branching fractions and the CP asymmetries for the considered quasi-two-body decay modes have been calculated by using the quasi-two-body approximation. It is found that the branching fractions are within the range of $10^{-9}-10^{-6}$. Our predicted branching fraction for $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$ agrees with the upper limit issued by the BABAR collaboration but is smaller than the LHCb measurement by a factor of ∼$\sim 2$. In Figs. (2, 3), we have shown the differential branching fraction over the $\pi$-$\pi$ invariant mass and the direct CP violation for the channel $B^{+}\to\pi^{+}f_{0}(500)\to\pi^{+}\pi^{+}\pi^{-}$. We hope all the predictions could be tested by further experiments, it may deeper our understanding for the $B$ meson decay mechanism in the context of PQCD factorization approach, and help us to probe the inner structure of $f_{0}(500)$.
Acknowledgements.
This work is supported by National Natural Science Foundation of China under Contract No.12275036, No.12347101, No.12265010, the Natural Science Foundation of Chongqing under Contract No.cstc2021jcyjmsxmX0681, the Science and Technology Research Program of Chongqing Municipal Education Commission under Contract No.KJQN202001541 and the Research Foundation of Chongqing University of Science and Technology under the project No.ckrc20231220.
Appendix A Decay Amplitudes
$F^{LL}_{T\sigma}$ $=$ $8\pi C_{F}m^{4}_{B}f_{P}\int dx_{B}dz\int b_{B}db_{B}bdb\phi_{B}(x_{B},b_{B})(1-\eta)\big{\{}\big{[}\sqrt{\eta}(1-2z)(\phi^{s}+\phi^{t})+(1+z)\phi^{0}\big{]}$ (34)
$\times$ $E_{1ab}(t_{1a})h_{1a}(x_{B},z,b_{B},b)S_{t}(z)+\sqrt{\eta}\left(2\phi^{s}-\sqrt{\eta}\phi^{0}\right)E_{1ab}(t_{1b})h_{1b}(x_{B},z,b_{B},b)S_{t}(x_{B})\big{\}}\;,$
$F^{SP}_{T\sigma}$ $=$ −$-16\pi C_{F}m^{4}_{B}r_{0}f_{P}\int dx_{B}dz\int b_{B}db_{B}bdb\phi_{B}(x_{B},b_{B})\big{\{}\left[\sqrt{\eta}(2+z)\phi^{s}-\sqrt{\eta}z\phi^{t}+(1+\eta(1-2z))\phi^{0}\right]$ (35)
$\times$ $E_{1ab}(t_{1a})h_{1a}(x_{B},z,b_{B},b)S_{t}(z)+\left[2\sqrt{\eta}(1-x_{B}+\eta)\phi^{s}+(x_{B}-2\eta)\phi^{0}\right]$
$\times$ $E_{1ab}(t_{1b})h_{1b}(x_{B},z,b_{B},b)S_{t}(x_{B})\big{\}}\;,$
[…]
$\times$ $\big{\{}\left[\sqrt{\eta}z(\phi^{t}-\phi^{s})+((1-\eta)(1-x_{3})-x_{B}+z\eta)\phi^{0}\right]E_{1cd}(t_{1c})h_{1c}(x_{B},z,x_{3},b_{B},b_{3})$
$+$ $\left[z(\sqrt{\eta}(\phi^{s}+\phi^{t})-\phi^{0})-(x_{3}(1-\eta)-x_{B})\phi^{0}\right]E_{1cd}(t_{1d})h_{1d}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$M^{LR}_{T\sigma}$ $=$ −$-32\pi C_{F}r_{0}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}\sqrt{\eta}z(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})$ (37)
$+$ $\sqrt{\eta}((1-x_{3})(1-\eta)-x_{B})(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})+((1-x_{3})(1-\eta)-x_{B})(\phi^{P}+\phi^{T})\phi^{0}$
$+$ $\eta z(\phi^{P}-\phi^{T})\phi^{0}\big{]}E_{1cd}(t_{1c})h_{1c}(x_{B},z,x_{3},b_{B},b_{3})+\big{[}-\sqrt{\eta}z(\phi^{P}+\phi^{T})(\sqrt{\eta}\phi^{0}+(\phi^{t}+\phi^{s}))$
$+$ $(x_{B}-x_{3}(1-\eta))(\phi^{P}-\phi^{T})(\sqrt{\eta}(\phi^{s}-\phi^{t})+\phi^{0})\big{]}E_{1cd}(t_{1d})h_{1d}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
[…]
$\times$ $\big{\{}\left[\sqrt{\eta}z(\phi^{t}+\phi^{s})+((\eta-1)(1-x_{3})+x_{B}-z\eta)\phi^{0}\right]E_{1cd}(t_{1c})h_{1c}(x_{B},z,x_{3},b_{B},b_{3})$
$+$ $\left[z(\sqrt{\eta}(\phi^{s}-\phi^{t})-\eta\phi^{0})-(x_{3}(1-\eta)-x_{B})\phi^{0}\right]E_{1cd}(t_{1d})h_{1d}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$F^{LL}_{A\sigma}$ $=$ $8\pi C_{F}m^{4}_{B}f_{B}\int dzdx_{3}\int bdbb_{3}db_{3}\big{\{}\left[2r_{0}\sqrt{\eta}\phi^{P}((2-z)\phi^{s}+z\phi^{t})-(1-\eta)(1-z)\phi^{A}\phi^{0}\right]E_{1ef}(t_{1e})$ (39)
$\times$ $h_{1e}(z,x_{3},b,b_{3})S_{t}(z)+\big{[}2r_{0}\sqrt{\eta}[(1-x_{3})(1-\eta)\phi^{T}-(1+x_{3}+(1-x_{3})\eta)\phi^{P}]\phi^{s}$
$+$ $(x_{3}(1-\eta)+\eta)(1-\eta)\phi^{A}\phi^{0}\big{]}E_{1ef}(t_{1f})h_{1f}(z,x_{3},b,b_{3})S_{t}(x_{3})\big{\}}\;,$
$F^{SP}_{A\sigma}$ $=$ $16\pi C_{F}m^{4}_{B}f_{B}\int dzdx_{3}\int bdbb_{3}db_{3}\big{\{}\left[\sqrt{\eta}(1-\eta)(1-z)\phi^{A}(\phi^{s}+\phi^{t})-2r_{0}(1+(1-z)\eta)\phi^{P}\phi^{0}\right]$ (40)
$\times$ $E_{1ef}(t_{1e})h_{1e}(z,x_{3},b,b_{3})S_{t}(z)+\left[2\sqrt{\eta}(1-\eta)\phi^{A}\phi^{s}-r_{0}\left(2\eta\phi^{P}+x_{3}(1-\eta)(\phi^{P}-\phi^{T})\right)\phi^{0}\right]$
$\times$ $E_{1ef}(t_{1f})h_{1f}(z,x_{3},b,b_{3})S_{t}(x_{3})\big{\}}\;,$
$M^{LL}_{A\sigma}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}(\eta-1)[x_{3}(1-\eta)+x_{B}+\eta(1-z)]\phi^{A}\phi^{0}$ (41)
$+$ $r_{0}\sqrt{\eta}(x_{3}(1-\eta)+x_{B}+\eta)(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})+r_{0}\sqrt{\eta}(1-z)(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})$
$+$ $2r_{0}\sqrt{\eta}(\phi^{P}\phi^{s}+\phi^{T}\phi^{t})\big{]}E_{1gh}(t_{1g})h_{1g}(x_{B},z,x_{3},b_{B},b_{3})+\big{[}(1-\eta^{2})(1-z)\phi^{A}\phi^{0}$
$+$ $r_{0}\sqrt{\eta}(x_{B}-x_{3}(1-\eta)-\eta)(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})-r_{0}\sqrt{\eta}(1-z)(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})\big{]}$
$\times$ $E_{1gh}(t_{1h})h_{1h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$M^{LR}_{A\sigma}$ $=$ −$-32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}\sqrt{\eta}(1-\eta)(1+z)\phi^{A}(\phi^{s}-\phi^{t})$ (42)
$+$ $r_{0}(2-x_{B}-x_{3}(1-\eta))(\phi^{P}+\phi^{T})\phi^{0}+r_{0}\eta(z\phi^{P}-(2+z)\phi^{T})\phi^{0}\big{]}E_{1gh}(t_{1g})h_{1g}(x_{B},z,x_{3},b_{B},b_{3})$
$+$ $\big{[}\sqrt{\eta}(1-\eta)(1-z)\phi^{A}(\phi^{s}-\phi^{t})+r_{0}((x_{3}(1-\eta)-x_{B})(\phi^{P}+\phi^{T})+\eta((2-z)\phi^{P}+z\phi^{T}))\phi^{0}\big{]}$
$\times$ $E_{1gh}(t_{1h})h_{1h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$M^{SP}_{A\sigma}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}\phi^{0}\phi^{A}(\eta-1)(z(1+\eta)-1)$ (43)
$+$ $r_{0}\sqrt{\eta}(\phi^{T}-\phi^{P})(\phi^{s}+\phi^{t})((\eta-1)(1-x_{3})+x_{B})+r_{0}\sqrt{\eta}(\phi^{T}+\phi^{P})(\phi^{s}-\phi^{t})z-4r_{0}\sqrt{\eta}\phi^{P}\phi^{s}\big{]}$
$\times$ $E_{1gh}(t_{1g})h_{1g}(x_{B},z,x_{3},b_{B},b_{3})+\big{[}\phi^{0}\phi^{A}(\eta-1)(x3(1-\eta)+\eta(2-z)-x_{B})$
$+$ $(\phi^{T}+\phi^{P})(\phi^{s}-\phi^{t})r_{0}\sqrt{\eta}(x_{3}(1-\eta)+\eta-x_{B})+(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})r_{0}\sqrt{\eta}(1-z)\big{]}$
$\times$ $E_{1gh}(t_{1h})h_{1h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$F^{SP}_{TP}$ $=$ $16\pi C_{F}F_{s}(\omega^{2})\sqrt{\eta}m^{4}_{B}\int dx_{B}dx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}(\eta-1)\phi^{A}+r_{0}x_{3}(\eta-1)(\phi^{P}-\phi^{T})$ (44)
-- $2r_{0}\phi^{P}\big{]}E_{2ab}(t_{2a})h_{2a}(x_{B},x_{3},b_{B},b_{3})S_{t}(x_{3})+\left[x_{B}(\eta-1)\phi^{A}+2r_{0}(\eta+x_{B}-1)\phi^{P}\right]$
$\times$ $E_{2ab}(t_{2b})h_{2b}(x_{B},x_{3},b_{B},b_{3})S_{t}(x_{B})\big{\}}\;,$
$M^{LL}_{TP}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}bdb\phi_{B}(x_{B},b_{B})\phi^{0}\big{\{}\big{[}(1-x_{B}-z)(1-\eta^{2})\phi^{A}$ (45)
-- $r_{0}x_{3}(1-\eta)(\phi^{P}-\phi^{T})+r_{0}(x_{B}+z)\eta(\phi^{P}+\phi^{T})-2r_{0}\eta\phi^{P}\big{]}E_{2cd}(t_{2c})h_{2c}(x_{B},z,x_{3},b_{B},b)$
-- $\big{[}(z-x_{B}+x_{3}(1-\eta))(1-\eta)\phi^{A}+r_{0}(x_{B}-z)\eta(\phi^{P}-\phi^{T})-r_{0}x_{3}(1-\eta)(\phi^{P}+\phi^{T})\big{]}$
$\times$ $E_{2cd}(t_{2d})h_{2d}(x_{B},z,x_{3},b_{B},b)\big{\}}\;,$
$M^{LR}_{TP}$ $=$ $32\pi C_{F}m^{4}_{B}\sqrt{\eta}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}bdb\phi_{B}(x_{B},b_{B})\big{\{}\big{[}(1-x_{B}-z)(1-\eta)(\phi^{s}+\phi^{t})\phi^{A}$ (46)
$+$ $r_{0}(1-x_{B}-z)(\phi^{s}+\phi^{t})(\phi^{P}-\phi^{T})+r_{0}(x_{3}(1-\eta)+\eta)(\phi^{s}-\phi^{t})(\phi^{P}+\phi^{T})\big{]}$
$\times$ $E_{2cd}(t_{2c})h_{2c}(x_{B},z,x_{3},b_{B},b)-\big{[}(z-x_{B})(1-\eta)(\phi^{s}-\phi^{t})\phi^{A}+r_{0}(z-x_{B})(\phi^{s}-\phi^{t})(\phi^{P}-\phi^{T})$
$+$ $r_{0}x_{3}(1-\eta)(\phi^{s}+\phi^{t})(\phi^{P}+\phi^{T})\big{]}E_{2cd}(t_{2d})h_{2d}(x_{B},z,x_{3},b_{B},b)\big{\}}\;,$
$M^{SP}_{TP}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}bdb\phi_{B}(x_{B},b_{B})\phi^{0}\big{\{}\big{[}(1+\eta-x_{B}-z+x_{3}(1-\eta))(1-\eta)\phi^{A}$ (47)
$+$ $r_{0}\eta(x_{B}+z)(\phi^{P}-\phi^{T})-r_{0}x_{3}(1-\eta)(\phi^{P}+\phi^{T})-2r_{0}\eta\phi^{P}\big{]}E_{2cd}(t_{2c})h_{2c}(x_{B},z,x_{3},b_{B},b)$
-- $\left[(z-x_{B})(1-\eta^{2})\phi^{A}-r_{0}x_{3}(1-\eta)(\phi^{P}-\phi^{T})+r_{0}\eta(x_{B}-z)(\phi^{P}+\phi^{T})\right]$
$\times$ $E_{2cd}(t_{2d})h_{2d}(x_{B},z,x_{3},b_{B},b)\big{\}}\;,$
$F^{LL}_{AP}$ $=$ $8\pi C_{F}m^{4}_{B}f_{B}\int dzdx_{3}\int bdbb_{3}db_{3}\big{\{}\big{[}(x_{3}(1-\eta)-1)(1-\eta)\phi^{A}\phi^{0}+2r_{0}\sqrt{\eta}(x_{3}(1-\eta)(\phi^{P}-\phi^{T})$ (48)
-- $2\phi^{P})\phi^{s}\big{]}E_{2ef}(t_{2e})h_{2e}(z,x_{3},b,b_{3})S_{t}(x_{3})+\big{[}z(1-\eta)\phi^{A}\phi^{0}+2r_{0}\sqrt{\eta}\phi^{P}((1-\eta)(\phi^{s}-\phi^{t})$
$+$ $z(\phi^{s}+\phi^{t}))\big{]}E_{2ef}(t_{2f})h_{2f}(z,x_{3},b,b_{3})S_{t}(z)\big{\}}\;,$
$F^{SP}_{AP}$ $=$ $16\pi C_{F}m^{4}_{B}f_{B}\int dzdx_{3}\int bdbb_{3}db_{3}\big{\{}\big{[}2\sqrt{\eta}(1-\eta)\phi^{A}\phi^{s}+r_{0}(1-x_{3})(\phi^{P}+\phi^{T})\phi^{0}$ (49)
$+$ $r_{0}\eta((1+x_{3})\phi^{P}-(1-x_{3})\phi^{T})\phi^{0}\big{]}E_{2ef}(t_{2e})h_{2e}(z,x_{3},b,b_{3})S_{t}(x_{3})$
$+$ $\left[2r_{0}(1-\eta(1-z))\phi^{P}\phi^{0}+z\sqrt{\eta}((1-\eta)\phi^{A}(\phi^{s}-\phi^{t})\right]E_{2ef}(t_{2f})h_{2f}(z,x_{3},b,b_{3})S_{t}(z)\big{\}}\;,$
$M^{LL}_{AP}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}(\eta-1)(-\eta+(1+\eta)(x_{B}+z))\phi^{A}\phi^{0}$ (50)
$+$ $r_{0}\sqrt{\eta}(x_{3}(1-\eta)+\eta)(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})+r_{0}\sqrt{\eta}(1-x_{B}-z)(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})-4r_{0}\sqrt{\eta}\phi^{P}\phi^{s}\big{]}$
$\times$ $E_{2gh}(t_{2g})h_{2g}(x_{B},z,x_{3},b_{B},b_{3})+\big{[}(1-\eta)((1-x_{3})(1-\eta)-\eta(x_{B}-z))\phi^{A}\phi^{0}-r_{0}\sqrt{\eta}(x_{B}-z)$
$\times$ $(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})+r_{0}\sqrt{\eta}(1-\eta)(1-x_{3})(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})\big{]}E_{2gh}(t_{2h})h_{2h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$M^{LR}_{AP}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}\sqrt{\eta}(1-\eta)(2-x_{B}-z)\phi^{A}(\phi^{s}+\phi^{t})$ (51)
-- $r_{0}(1+x_{3})(\phi^{P}-\phi^{T})\phi^{0}-r_{0}\eta[(1-x_{B}-z)(\phi^{P}+\phi^{T})-x_{3}(\phi^{P}-\phi^{T})+2\phi^{P}]\phi^{0}\big{]}$
$\times$ $E_{2gh}(t_{2g})h_{2g}(x_{B},z,x_{3},b_{B},b_{3})-\big{[}r_{0}(1-\eta)(x_{3}-1)(\phi^{P}-\phi^{T})\phi^{0}+\sqrt{\eta}(x_{B}-z)[r_{0}\sqrt{\eta}(\phi^{P}+\phi^{T})\phi^{0}$
$+$ $(1-\eta)\phi^{A}(\phi^{s}+\phi^{t})]\big{]}E_{2gh}(t_{2h})h_{2h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;,$
$M^{SP}_{AP}$ $=$ $32\pi C_{F}m^{4}_{B}/\sqrt{2N_{c}}\int dx_{B}dzdx_{3}\int b_{B}db_{B}b_{3}db_{3}\phi_{B}(x_{B},b_{B})\big{\{}\big{[}(\eta-1)\phi^{0}\phi^{A}(\eta(x_{3}+x_{B}+z-2)$ (52)
-- $x_{3}+1)+r_{0}\sqrt{\eta}[(1-\eta)(1-x_{3})(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})+(z+x_{B})(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})$
$+$ $2(\phi^{P}\phi^{s}+\phi^{T}\phi^{t})]\big{]}E_{2gh}(t_{2g})h_{2g}(x_{B},z,x_{3},b_{B},b_{3})-\big{[}r_{0}\sqrt{\eta}(\phi^{P}+\phi^{T})(\phi^{s}-\phi^{t})(1-\eta)(x_{3}-1)$
$+$ $r_{0}\sqrt{\eta}(\phi^{P}-\phi^{T})(\phi^{s}+\phi^{t})(x_{B}-z)+\phi^{0}\phi^{A}(1-\eta^{2})(x_{B}-z)\big{]}E_{2gh}(t_{2h})h_{2h}(x_{B},z,x_{3},b_{B},b_{3})\big{\}}\;.$
where the hard functions are defined as
$h_{i}(x_{1},x_{2},(x_{3},)b_{1},b_{2})$ $=$ $h_{1}(\beta,b_{2})\times h_{2}(\alpha,b1,b_{2})$
$h_{1}(\beta,b_{2})$ $=$ $\left\{\begin{array}[]{ll}K_{0}(\sqrt{\beta}b_{2}),&\beta\geq 0,\\ \frac{i\pi}{2}H^{(1)}_{0}(\sqrt{-\beta}b_{2}),&\beta<0,\end{array}\right.$ (55)
$h_{2}(\alpha,b1,b_{2})$ $=$ $\left\{\begin{array}[]{ll}\theta(b_{2}-b_{1})K_{0}(\sqrt{\alpha}b_{2})I_{0}(\sqrt{\alpha}b_{1}),&\alpha\geq 0,\\ \theta(b_{2}-b_{1})\frac{i\pi}{2}H^{(1)}_{0}(\sqrt{-\alpha}b_{2})J_{0}(\sqrt{-\alpha}b_{1}),&\alpha<0,\end{array}\right.$ (58)
where $E_{1mn},E_{2mn},E_{3mn},E_{4mn}$($m=a,c,e,g$ and $n=b,d,f,h$) are the evolution factors,
$E_{1ab}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)],$
$E_{1cd}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)-S_{P}(t)]_{b=b_{B}},$
$E_{1ef}(t)$ $=$ $\alpha(t)\exp[-S_{P}(t)-S_{\sigma}(t)],$
$E_{1gh}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)-S_{P}(t)]_{b=b_{3}},$
$E_{2ab}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{P}(t)],$
$E_{2cd}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)-S_{P}(t)]_{b_{3}=b_{B}},$
$E_{2ef}(t)$ $=$ $E_{1ef}(t),$
$E_{2gh}(t)$ $=$ $E_{1gh}(t),$
$E_{3ab}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)],$
$E_{3cd}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)-S_{D}(t)]_{b=b_{B}},$
$E_{3ef}(t)$ $=$ $\alpha(t)\exp[-S_{D}(t)-S_{\sigma}(t)],$
$E_{3gh}(t)$ $=$ $\alpha(t)\exp[-S_{B}(t)-S_{\sigma}(t)-S_{D}(t)]_{b=b_{3}},$
$E_{4ef}(t)$ $=$ $E_{3ef}(t),$
$E_{4gh}(t)$ $=$ $E_{3gh}(t)\,.$ (59)
The Sudakov form factors are defined as
$S_{B}(t)=s\left(\frac{x_{B}m_{B}}{\sqrt{2}},b_{B}\right)+\frac{5}{3}\int^{t}_{1/b_{B}}\frac{d\bar{\mu}}{\bar{\mu}}\gamma_{q}(\alpha_{s}(\bar{\mu})),$
$S_{\sigma}(t)=s\left(\frac{z(1-r^{2})m_{B}}{\sqrt{2}},b\right)+s\left(\frac{(1-z)(1-r^{2})m_{B}}{\sqrt{2}},b\right)+2\int^{t}_{1/b}\frac{d\bar{\mu}}{\bar{\mu}}\gamma_{q}(\alpha_{s}(\bar{\mu})),$
$S_{D,P}(t)=s\left(\frac{x_{3}m_{B}}{\sqrt{2}},b_{3}\right)+s\left(\frac{(1-x_{3})m_{B}}{\sqrt{2}},b_{3}\right)+2\int^{t}_{1/b_{3}}\frac{d\bar{\mu}}{\bar{\mu}}\gamma_{q}(\alpha_{s}(\bar{\mu}))\,.$ (60)
Practically, the mentioned hard scales are chosen as
$t_{1a}={\rm Max}\left\{\sqrt{|\alpha_{1a}|},\sqrt{|\beta_{1a}|},1/b_{B},1/b\right\},$ $t_{1b}={\rm Max}\left\{\sqrt{|\alpha_{1b}|},\sqrt{|\beta_{1b}|},1/b_{B},1/b\right\},$
$t_{1c}={\rm Max}\left\{\sqrt{|\alpha_{1c}|},\sqrt{|\beta_{1c}|},1/b_{B},1/b_{3}\right\},$ $t_{1d}={\rm Max}\left\{\sqrt{|\alpha_{1d}|},\sqrt{|\beta_{1d}|},1/b_{B},1/b_{3}\right\},$
$t_{1e}={\rm Max}\left\{\sqrt{|\alpha_{1e}|},\sqrt{|\beta_{1e}|},1/b_{3},1/b\right\},$ $t_{1f}={\rm Max}\left\{\sqrt{|\alpha_{1f}|},\sqrt{|\beta_{1f}|},1/b_{3},1/b\right\},$
$t_{1g}={\rm Max}\left\{\sqrt{|\alpha_{1g}|},\sqrt{|\beta_{1g}|},1/b_{B},1/b_{3}\right\},$ $t_{1h}={\rm Max}\left\{\sqrt{|\alpha_{1h}|},\sqrt{|\beta_{1h}|},1/b_{B},1/b_{3}\right\},$
$t_{2a}={\rm Max}\left\{\sqrt{|\alpha_{2a}|},\sqrt{|\beta_{2a}|},1/b_{B},1/b_{3}\right\},$ $t_{2b}={\rm Max}\left\{\sqrt{|\alpha_{2b}|},\sqrt{|\beta_{2b}|},1/b_{B},1/b_{3}\right\},$
$t_{2c}={\rm Max}\left\{\sqrt{|\alpha_{2c}|},\sqrt{|\beta_{2c}|},1/b_{B},1/b\right\},$ $t_{2d}={\rm Max}\left\{\sqrt{|\alpha_{2d}|},\sqrt{|\beta_{2d}|},1/b_{B},1/b\right\},$
$t_{2e}={\rm Max}\left\{\sqrt{|\alpha_{2e}|},\sqrt{|\beta_{2e}|},1/b_{3},1/b\right\},$ $t_{2f}={\rm Max}\left\{\sqrt{|\alpha_{2f}|},\sqrt{|\beta_{2f}|},1/b_{3},1/b\right\},$
$t_{2g}={\rm Max}\left\{\sqrt{|\alpha_{2g}|},\sqrt{|\beta_{2g}|},1/b_{B},1/b_{3}\right\},$ $t_{2h}={\rm Max}\left\{\sqrt{|\alpha_{2h}|},\sqrt{|\beta_{2h}|},1/b_{B},1/b_{3}\right\},$
$t_{3a}={\rm Max}\left\{\sqrt{|\alpha_{3a}|},\sqrt{|\beta_{3a}|},1/b_{B},1/b\right\},$ $t_{3b}={\rm Max}\left\{\sqrt{|\alpha_{3b}|},\sqrt{|\beta_{3b}|},1/b_{B},1/b\right\},$
$t_{3c}={\rm Max}\left\{\sqrt{|\alpha_{3c}|},\sqrt{|\beta_{3c}|},1/b_{B},1/b_{3}\right\},$ $t_{3c}^{\prime}={\rm Max}\left\{\sqrt{|\alpha_{3c}|},\sqrt{|\beta^{\prime}_{1c}|},1/b_{B},1/b_{3}\right\},$
$t_{3d}={\rm Max}\left\{\sqrt{|\alpha_{3d}|},\sqrt{|\beta_{3d}|},1/b_{B},1/b_{3}\right\},$ $t_{3d}^{\prime}={\rm Max}\left\{\sqrt{|\alpha_{3d}|},\sqrt{|\beta^{\prime}_{1d}|},1/b_{B},1/b_{3}\right\},$
$t_{4e}={\rm Max}\left\{\sqrt{|\alpha_{4e}|},\sqrt{|\beta_{4e}|},1/b_{3},1/b\right\},$ $t_{4f}={\rm Max}\left\{\sqrt{|\alpha_{4f}|},\sqrt{|\beta_{4f}|},1/b_{3},1/b\right\},$
$t_{4g}={\rm Max}\left\{\sqrt{|\alpha_{4g}|},\sqrt{|\beta_{4g}|},1/b_{B},1/b_{3}\right\},$ $t_{4h}={\rm Max}\left\{\sqrt{|\alpha_{4h}|},\sqrt{|\beta_{4h}|},1/b_{B},1/b_{3}\right\}.$ (61)
The jet function $S_{t}$ resums the threshold double logarithm and can be parameterized as
$\frac{2^{1+2c}\Gamma(3/2+c)}{\sqrt{\pi}\Gamma(1+c)}[x(1-x)]^{c}\;,$ (62)
with $c=0.4$ for numerical calculation.
$\alpha_{1a}$ $=$ $zm_{B}^{2}$
$\beta_{1a}$ $=$ $x_{B}zm_{B}^{2}=\beta_{1b}=\alpha_{1c}=\alpha_{1d}$
$\alpha_{1b}$ $=$ $(x_{B}-\eta)m_{B}^{2}$
$\beta_{1c}$ $=$ −$-z[(1-\eta)(1-x_{3})-x_{B}]m_{B}^{2}$
$\beta_{1d}$ $=$ −$-z[(1-\eta)x_{3}-x_{B}]m_{B}^{2}$
$\alpha_{1e}$ $=$ $(z-1)m_{B}^{2}$
$\beta_{1e}$ $=$ −$-(1-z)[\eta+(1-\eta)x_{3}]m_{B}^{2}=\beta_{1f}=\alpha_{1g}=\alpha_{1h}$
$\alpha_{1f}$ $=$ −$-(\eta+(1-\eta)x_{3})m_{B}^{2}$
$\beta_{1g}$ $=$ $\{1-z[(1-\eta)(1-x_{3})-x_{B}]\}m_{B}^{2}$
$\beta_{1h}$ $=$ −$-(1-z)[(1-\eta)x_{3}+\eta-x_{B}]m_{B}^{2}$
$\alpha_{2a}$ $=$ $x_{3}(1-\eta)m_{B}^{2}$
$\beta_{2a}$ $=$ $x_{3}x_{B}(1-\eta)m_{B}^{2}=\beta_{2b}=\alpha_{2c}=\alpha_{2d}$
$\alpha_{2b}$ $=$ $x_{B}(1-\eta)m_{B}^{2}$
$\beta_{2c}$ $=$ $(1-x_{B}-z)[(\eta-1)x_{3}-\eta]m_{B}^{2}$
$\beta_{2d}$ $=$ $(1-\eta)(x_{B}-z)m_{B}^{2}$
$\alpha_{2e}$ $=$ −$-[1-x_{3}(1-\eta)]m_{B}^{2}$
$\beta_{2e}$ $=$ −$-z(1-\eta)(1-x_{3})m_{B}^{2}=\beta_{2f}=\alpha_{2g}=\alpha_{2h}$
$\alpha_{2f}$ $=$ $z(\eta-1)m_{B}^{2}$
$\beta_{2g}$ $=$ $\{1-(1-z-x_{B})[\eta+(1-\eta)x_{3}]\}m_{B}^{2}$
$\beta_{2h}$ $=$ $(x_{B}-z)(1-\eta)(1-x_{3})m_{B}^{2}$
$\alpha_{3a}$ $=$ $z(1-r^{2})m_{B}^{2}$
$\beta_{3a}$ $=$ $x_{B}z(1-r^{2})m_{B}^{2}=\beta_{3b}=\alpha_{3c}=\alpha_{3d}=\alpha^{\prime}_{3c}=\alpha^{\prime}_{3d}$
$\alpha_{3b}$ $=$ $(1-r^{2})(x_{B}-\eta)m_{B}^{2}$
$\beta_{3c}$ $=$ −$-[z(1-r^{2})+r^{2}(1-x_{3})][(1-\eta)(1-x_{3})-x_{B}]m_{B}^{2}$
$\beta_{3d}$ $=$ $\{r_{c}^{2}-[z(1-r^{2})+x_{3}r^{2}][(1-\eta)x_{3}-x_{B}]\}m_{B}^{2}$
$\beta^{\prime}_{3c}$ $=$ $\{r_{c}^{2}-[z(1-r^{2})+r^{2}(1-x_{3})][(1-\eta)(1-x_{3})-x_{B}]\}m_{B}^{2}$
$\beta^{\prime}_{3d}$ $=$ −$-[z(1-r^{2})+x_{3}r^{2}][(1-\eta)x_{3}-x_{B}]m_{B}^{2}$
$\alpha_{3e}$ $=$ −$-[1-z(1-r^{2})-r_{c}^{2}]m_{B}^{2}$
$\beta_{3e}$ $=$ −$-[(1-r^{2})(1-z)+x_{3}r^{2}][\eta+(1-\eta)x_{3}]m_{B}^{2}=\beta_{3f}=\alpha_{3g}=\alpha_{3h}$
$\alpha_{3f}$ $=$ −$-(1-r^{2}+x_{3}r^{2})(\eta+(1-\eta)x_{3})m_{B}^{2}$
$\beta_{3g}$ $=$ $\{1-[z(1-r^{2})+r^{2}(1-x_{3})][(1-\eta)(1-x_{3})-x_{B}]\}m_{B}^{2}$
$\beta_{3h}$ $=$ −$-[(1-z)(1-r^{2})+r^{2}x_{3}][(1-\eta)x_{3}+\eta-x_{B}]m_{B}^{2}$
$\alpha_{4e}$ $=$ −$-(1-x_{3}r^{2})[1-x_{3}(1-\eta)]m_{B}^{2}$
$\beta_{4e}$ $=$ −$-[(1-x_{3})r^{2}+z(1-r^{2})](1-\eta)(1-x_{3})m_{B}^{2}=\beta_{4f}=\alpha_{4g}=\alpha_{4h}$
$\alpha_{4f}$ $=$ $\{r_{c}^{2}-[r^{2}+z(1-r^{2})](1-\eta)\}m_{B}^{2}$
$\beta_{4g}$ $=$ $\{1-[(1-r^{2})(1-z)+x_{3}r^{2}-x_{B}][\eta+(1-\eta)x_{3}]\}m_{B}^{2}$
$\beta_{4h}$ $=$ −$-[r^{2}(1-x_{3})+z(1-r^{2})-x_{B}](1-\eta)(1-x_{3})m_{B}^{2}$ (63)